Circular No. 6290

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

GRO J1744−28

M. S. Briggs, University of Alabama, Huntsville (UAH); B. A. Harmon, NASA Marshall Space Flight Center (MSFC); J. van Paradijs, UAH and University of Amsterdam; C. Kouveliotou, Universities Space Research Association; G. J. Fishman, MSFC; and J. Kommers and W. H. G. Lewin, Massachusetts Institute of Technology, report for the Compton Observatory BATSE team: “The increased intensity of GRO J1744−28 enables improved spectral analysis. Using burst observations with the BATSE Large Area Detectors (> 30 keV) and a Spectroscopy Detector (> 8 keV), unabsorbed power law and OTTB models may be excluded. Acceptable fits are obtained with a blackbody model with no absorption and kT = 6 keV and with an OTTB model with kT = 10 keV and absorption of N(H) = 1.6 × 10^{24} cm^{−2}. Very similar spectral results are obtained for the persistent spectrum from occultation observations. This similarity argues for a common emission process for the burst and persistent emission.”

Fishman, Harmon, Kouveliotou, van Paradijs, Briggs, Kommers and Lewin, with also K. Deal and P. Woods, UAH, report: “The unusual burster in the galactic center region, identified with the x-ray pulsar GRO J1744−28 (IAUC 6286), continues to increase in burst intensity and steady emission. Over the past four weeks, both the peak flux and the fluence of the individual bursts have increased by about a factor of six. On Jan. 15 their average values (20–50 keV) were ∼ 2 and ∼ 7 × 10^{−7} erg cm^{−2} s^{−1}, respectively. Over the past ten days, the burstrate has increased to ∼ 25 observed bursts per day (∼ 40 per day, corrected for earth occultation and livetime) from a previous average of ∼ 18 observed bursts per day during the preceding month. The burst durations have remained the same, 4–6 s FWHM, over the entire period of observation. The associated steady source, as observed by earth occultation, has increased in intensity to ∼ 2.5 Crab (20–100 keV) on Jan. 16.”

COMET 6P/D’ARREST Total visual magnitudes: 1995 Aug. 25.56 UT, 8.0, (P. Camilleri, Co-

bram, Vic., Australia, 20 × 80 binoculars); Sept. 1.34, 7.8 (A. Hale, Cloudcroft, NM, 10 × 50 binoculars); 4.40, 7.2 (C. S. Morris, southern California, 10 × 50 binoculars); 17.25, 8.4 (Hale); 24.35, 8.2 (Morris); 30.04, 7.8 (J. Carvajal, Madrid, Spain, 10 × 70 binoculars); Oct. 3.41, 8.9 (Hale); 15.23, 9.7 (Hale, 0.41-m reflector); 21.28, 9.3 (Morris, 20 × 80 binoculars); 27.21, 10.2 (Hale); Nov. 14.18, 11.0 (Hale); 20.26, 11.3 (Hale); 25.24, 11.4 (Hale).

1996 January 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Brian G. Marsden